OpenStudy (anonymous):

4. It takes 7.8 kJ to raise the temperature of a 2.78 kg sample of an unknown substance by 12 C. Calculate its specific heat

OpenStudy (abmon98):

E=mcT E is the energy transferred in joules, J m is the mass of the substances in kg c is the specific heat capacity in J / kg °C T is the temperature change 7.8=2.78*12*C C=7.8/2.78*12

OpenStudy (anonymous):

Thank you, that's what I came up with also!

OpenStudy (abmon98):

sounds good your most welcome :)